OpenStudy (jungleman03):

Given that f(x) = 2x + 1 and g(x) = −5x + 2, solve for f(g(x)) when x = 3

OpenStudy (anonymous):

so 2(3) IS 6+1 = AND -5(3) IS -15+2=

OpenStudy (anonymous):

\[f(g(x)) = 2(-5x+2) + 1\] \[f(g(x))-> x=3: 2(-5(3) + 2) + 1 = -25.\]

OpenStudy (jungleman03):

thank you

OpenStudy (anonymous):

u r not suposed to give awnsers

OpenStudy (anonymous):

@green1